A flexible container at an initial volume of 3.10 L contains 8.51 mol of gas. More gas is then added to the container until it reaches a final volume of 13.7 L. Assuming the pressure and temperature of the gas remain constant, calculate the number of moles of gas added to the container.


1
Expert's answer
2021-11-10T00:30:27-0500

Avogadro's Law: "Equal volumes contain equal number of moles if pressure and temperature of the gas remain constant". V1/n1=V2/n2, where V is the volume and n is the number of moles.

3.10 L / 8.51 mol = 13.7 L / x mol

x = 8.51 mol * 13.7 L / 3.10 L = 37.61 mol

So, 13.7 L, n = 37.61 moles.

Since we started with 8.51 moles,

37.61 - 8.51= 29.1 moles added